# The Size-Mass relation at Rest-Frame m from JWST/NIRCam in the COSMOS-WEB and PRIMER-COSMOS fields

### v1

@@ 3.3 Profile Fits with GalfitM
The convolution box used by GalfitM is set to be 200 pixels per side. Despite slowing down the computation, the large convolution box was adopted to encompass even the largest sources. We can then use the same convolution box for all the targets avoiding the introduction of systematics due to the use of different convolution boxes for different galaxies.
Sérsic profile fits are independently fitted to the F277W and F444W images, allowing the Sérsic index to span the range 0.2 - 10 with an initial guess of 2.3. The x and y coordinates of the galaxies’ centers are limited to lie within a distance of 5 pixels from the SEP-based centers. The effective radius is limited to values between 0.01 and 200 pixels with the initial value defined as the semi-major axis of the SEP ellipse. Axis ratio and position angle are left as free parameters with no constraints. The mosaics are background-subtracted, but to allow for remaining background patterns across the mosaic we adopt a constant background with GalfitM as a free parameter.
@@
For 514 sources GalfitM produces a fit with . A visual inspection of these fits reveals that for the vast majority (506) the is driven by those galaxies fitted simultaneously with the target, and not by the target itself. We refit these sources, masking the nearby galaxies and we find that they converge properly, with much lower for 497 out of 506. The remainder are omitted from the final sample.
@@
Visual inspection by MM of all fitted galaxies shows 36 mergers and 2 lensing systems that are not properly fitted and, therefore, discarded (6 mergers and the 2 lensing systems were among those fits with ). We further reject 5 galaxies that were only partially covered by the NIRCam mosaics, leaving the fitting results unreliable. Our final sample with reliable Sérsic profiles includes galaxies of which are quiescent and are star-forming.
@@
In Figure 6 we show the GalfitM fitting results for three randomly selected galaxies. Although low redshift high-mass galaxies are bright and extended, our pipeline uses sufficiently large cutouts to enable accurate background estimates. At the same time, for low-mass high redshift galaxies we see that the SNR is sufficiently high for a precise size estimate (larger than 30 (50) for 84% (50%) of galaxies with and redshift ).
@@
The uncertainties retrieved for Sérsic index () and effective radius are systematically underestimated by GalfitM (van-der-wel12). Following the prescriptions in van-der-wel12 and martorano23, we reassess the uncertainties: we adopt a baseline uncertainty of 0.1 dex on the Sérsic index for galaxies with SNR = 50 and on for galaxies with SNR = 20, and compute the final uncertainties as . This methodology is based on an analysis of HST/WFC3 data (van-der-wel12). The SNR of JWST/NIRCam is much higher, producing, at times, very small uncertainties. In fact, for of the galaxies the GalfitM uncertainty on () is larger than the SNR-based value, and we retain the former.
@@
The rest-frame 1.5 and are calculated by fitting a straight line to the values of and in F277W and F444W, using the respective pivot wavelengths. This fit is repeated 100 times, sampling from the F277W and F444W and uncertainties, adopting the median values of the and polynomials at 1.5 as the estimates and half of the 16th-84th percentile range as the 1 uncertainty.

### v2

@@ 3.3 Profile Fits with GalfitM
The convolution box used by GalfitM is set to be 200 pixels per side. Despite slowing down the computation, the large convolution box was adopted to encompass even the largest sources. We can then use the same convolution box for all the targets, avoiding the introduction of systematics due to the use of different convolution boxes for different galaxies.
Sérsic profile fits are independently fitted to the F277W and F444W images, allowing the Sérsic index to span the range 0.2 - 10 with an initial guess of 2.3. The x and y coordinates of the galaxies’ centers are limited to lie within a distance of 5 pixels from the SEP-based centers. The effective radius is limited to values between 0.01 and 200 pixels with the initial value defined as the semi-major axis of the SEP ellipse. Axis ratio and position angle are left as free parameters with no constraints. The mosaics are background-subtracted, but to allow for remaining background patterns across the mosaic, we adopt a constant background with GalfitM as a free parameter.
@@
For 514 sources, GalfitM produces a fit with . A visual inspection of these fits reveals that for the vast majority (506) the is driven by those galaxies fitted simultaneously with the target and not by the target itself. We refit these sources, masking the nearby galaxies, and we find that they converge properly, with much lower for 497 out of 506. The remainder are omitted from the final sample.
@@
Visual inspection by M.M. of all fitted galaxies shows 36 mergers and two lensing systems that are not properly fitted and are, therefore, discarded (six mergers and the two lensing systems were among those fits with ). We further reject five galaxies that were only partially covered by the NIRCam mosaics, leaving the fitting results unreliable. Our final sample with reliable Sérsic profiles includes galaxies of which are quiescent and are star-forming.
@@
In Figure 6 we show the GalfitM fitting results for three randomly selected galaxies. Although low-redshift high-mass galaxies are bright and extended, our pipeline uses sufficiently large cutouts to enable accurate background estimates. At the same time, for low-mass high-redshift galaxies, we see that the signal-to-noise ratio (SNR) is sufficiently high for a precise size estimate (larger than 30 (50) for 84% (50%) of galaxies with and redshift ).
@@
The uncertainties retrieved for Sérsic index () and effective radius are systematically underestimated by GalfitM (van-der-wel12). Following the prescriptions in van-der-wel12 and martorano23, we reassess the uncertainties: we adopt a baseline uncertainty of 0.1 dex on the Sérsic index for galaxies with SNR = 50 and on for galaxies with SNR = 20, and compute the final uncertainties as . This methodology is based on an analysis of HST/WFC3 data (van-der-wel12). The SNR of JWST/NIRCam is much higher, producing, at times, very small uncertainties. In fact, for of the galaxies, the GalfitM uncertainty on () is larger than the SNR-based value, and we retain the former.
@@
The rest-frame 1.5 and are calculated by fitting a straight line to the values of and in F277W and F444W using the respective pivot wavelengths. This fit is repeated 100 times, sampling from the F277W and F444W and uncertainties, adopting the median values of the and polynomials at 1.5 as the estimates and half of the 16th-84th percentile range as the 1 uncertainty.
In Table 1 we present GalfitM outputs produced from the fitting of the NIRCam/F277W, F444W cutouts and the rest-frame effective radii and Sérsic indices computed as discussed above. A full version of the table is available online in the machine-readable format.
Label | Unit | Description |
Classic | — | COSMOS2020/CLASSIC Identifier [ID] |
RAdeg | deg | COSMOS2020/CLASSIC Right Ascension, decimal degrees (J2000) [ALPHA_J2000] |
DEdeg | deg | COSMOS2020/CLASSIC Declination, decimal degrees (J2000) [DELTA_2000] |
lpzBEST | — | COSMOS2020/CLASSIC LePhare photo-z redshift [lp_zBEST] |
loglpMassbest | [Msun] | Corrected (see text) COSMOS2020/CLASSIC LePhare BC03 log stellar mass [lp_mass_best] |
Re-F277W | arcsec | GalfitM effective radius in filter F277W |
e_Re-F277W | arcsec | Uncertainty in Re-F277W |
n-F277W | — | GalfitM Sérsic index in filter F277W |
e_n-F277W | — | Uncertainty in n-F277W |
mag-F277W | mag | GalfitM AB magnitude in filter F277W |
e_mag-F277W | mag | Uncertainty in mag-F277W |
q-F277W | — | GalfitM axis ratio in filter F277W |
e_q-F277W | — | Uncertainty in q-F277W |
PA-F277W | deg | GalfitM position angle in filter F277W |
e_PA-F277W | deg | Uncertainty in PA-F277W |
Re-F444W | arcsec | GalfitM effective radius in filter F444W |
e_Re-F444W | arcsec | Uncertainty in Re-F444W |
n-F444W | — | GalfitM Sérsic index in filter F444W |
e_n-F444W | — | Uncertainty in n-F444W |
mag-F444W | mag | GalfitM AB magnitude in filter F444W |
e_mag-F444W | mag | Uncertainty in mag-F444W |
q-F444W | — | GalfitM axis ratio in filter F444W |
e_q-F444W | — | Uncertainty in q-F444W |
PA-F444W | deg | GalfitM position angle in filter F444W |
e_PA-F444W | deg | Uncertainty in PA-F444W |
Re-1.5um | arcsec | Effective radius at rest-frame |
e_Re-1.5um | arcsec | Uncertainty in Re-1.5um |
n-1.5um | — | Sérsic index at rest-frame |
e_n-1.5um | — | Uncertainty in n-1.5um |
Note. — The
original COSMOS2020/CLASSIC labels are given in square brackets in the
description of the columns taken or derived from that catalog.
Table 1 is published in its entirety in
the machine-readable format. A description of the columns is provided here
for guidance regarding its form and content.